Physics 133 Challenging Problem 1 Two positively charged spheres of mass 1 kg each are held in free space, 1 meter away from each other. The two spheres have a charge of 1.0541 x 10-5C each. One charge is released while the other is still held fixed. a) Find the speed of the moving charge when it is 2 meters away from the fixed charge. b) Find the speed of the moving charge when it is 3 meters away from the fixed charge.
Answer #1

here,

charge , Q = 1.0541 * 10^-5 C

mass , m = 1 kg

r0 = 1 m

a)

r = 2 m

let the final speed of moving charge be v

using conservation of energy

K * q^2 * ( 1/r0 - 1/r) = 0.5 * m * v^2

9 * 10^9 * ( 1.0541 * 10^-5)^2 * ( 1 - 1/2) = 0.5 * 1 * v^2

solving for v

v = 1 m/s

b)

r = 3 m

let the final speed of moving charge be v

using conservation of energy

K * q^2 * ( 1/r0 - 1/r) = 0.5 * m * v^2

9 * 10^9 * ( 1.0541 * 10^-5)^2 * ( 1 - 1/3) = 0.5 * 1 * v^2

solving for v

v = 1.15 m/s